# Zhiwei Yang1,2    Chen Gao2†    Mike Zheng Shou2†

### v1

@@ 1 Introduction
Technically, the proposed PANDA is distinguished by the following aspects. (1) self-adaptive scene-aware strategy planning. Faced with a new scene or user-defined anomaly detection requirements, PANDA first conducts environment perception and understanding, then retrieves relevant anomaly rules from an anomaly knowledge database. Based on the environment context information, a scene-adaptive Retrieval Augmented Generation (RAG) mechanism is designed to construct tailored anomaly detection plans. (2) Goal-driven heuristic reasoning. PANDA injects task-specific prompts guided by latent anomaly cues, which steer the reasoning process toward more accurate and focused decision-making. (3) Tool-augmented self-reflection. PANDA iteratively assesses uncertainty and activates a suite of curated tools, such as object detection, image retrieve, or web search, to acquire additional information and resolve ambiguous decision-making. (4) Self-improving chain-of-memory. PANDA integrates historical experiences to justify current reasoning decisions or self-reflection. By progressively accumulating contextual cues across temporal spans, it enhances both the stability and accuracy of its decisions over time. Taken together, PANDA embodies an agentic AI engineer that proactively perceives diverse environments, formulates adaptive strategies, performs goal-driven reasoning, and progressively improves through tool-augmented reflection and a chain-of-memory mechanism, enabling robust and generalizable video anomaly detection across complex real-world scenarios.
@@ 2 Related Work
These approaches rely on varying levels of annotated data and typically fall into three categories: semi-supervised VAD hasan2016learning-12 ; lu2013abnormal-18 ; liu2018future-13 ; yang2023video-17 ; yang2022dynamic-15 , weakly-supervised VAD yang2024text-21 ; wu2024vadclip-25 ; wu2020not-24 ; sultani2018real-22 ; wu2024weakly , and instruction-tuned VAD du2024uncovering-29 ; tang2024hawk-30 ; zhang2024holmes-07 ; zhang2024holmes-32 . For example, Ristea et al. ristea2024self proposed an efficient anomaly detection model based on a lightweight masked autoencoder. Yang et al. yang2024text-21 introduced a text prompt-driven pseudo-labeling and self-training framework for weakly-supervised VAD. Zhang et al. zhang2024holmes-32 presented a model combining an anomaly-focused temporal sampler with an instruction-tuned MLLM to detect anomalies. While these training-dependent methods often perform well within the domain of the training data, they typically suffer from sharp performance degradation when deployed in out-of-distribution environments or faced with novel anomaly types. This limits their applicability in the open-world scenarios where anomalies are diverse, and context-sensitive.
@@ 3.1 Self-adaptive Scene-aware Strategy Planning
Given a user-defined detection query $\texttt{User}_{\text{query}}$ and an input video sequence $V=\{f_{1},f_{2},\dots,f_{N}\}$ containing $N$ frames, PANDA uniformly samples $M$ keyframes $F=\{f_{1},f_{2},\dots,f_{M}\}$ and constructs a perception prompt combining $F$ with the $\texttt{User}_{\text{query}}$. This prompt is fed to a VLM, which returns structured environmental information EnvInfo = {Scene Overview, Potential Anomalies, Weather Condition, Video Quality}. Here, Scene Overview provides a high-level summary of the scene, including location type (e.g., street, shop, parking lot) and observed activities. Potential Anomalies refers to types of suspicious behaviors that may plausibly occur in the current scene context. Weather Condition captures attributes such as time of day (day/night) and weather (e.g., sunny, rainy). Video Quality summarizes resolution and clarity (e.g., low-resolution, blurred, noisy).
@@ 3.2 Goal-Driven Heuristic Reasoning
Under the guidance of the detection strategy plan constructed in subsection 3.1, PANDA performs goal-driven heuristic reasoning using a VLM. Given the $\texttt{User}_{\text{query}}$, a clip-level video segment $V_{\text{clip}}=\{c_{1},c_{2},\dots,c_{T}\}$, here each video clip $c_{i}$ contains $s$ video frames, and the $\texttt{Plan}_{\texttt{strategy}}$, PANDA first applies the preprocessing tools specified in $\texttt{Plan}_{\texttt{strategy}}$ to obtain an enhanced video clip:
@@
Promptreasoning = $\text{Memory}^{\text{$l$-steps}}_{\text{text}}$, Potential Anomalies, Anomaly Detection Rules, Heuristic Prompt, Enhancement and Reflection Info}.
The fields Potential Anomalies, Anomaly Detection Rules, and Heuristic Prompt are directly inherited from the planning stage. The Enhancement and Reflection Info field incorporates information produced during the self-reflection stage (To be described in subsection 3.3), including tool-based refinements and updated anomaly rules and heuristic prompts. To enhance temporal awareness, PANDA equips a short-term memory component $\texttt{Memory}^{l\text{-steps}}_{\text{text}}$, which records the past $l$ reasoning steps as textual memory. In addition to textual memory, PANDA also maintains a corresponding visual memory stream $\texttt{Memory}^{l\text{-steps}}_{\text{visual}}$, which stores visual frames aligned with the latest $l$ steps, allowing the model to access fine-grained visual cues during inference.
@@
$\texttt{Result}_{\text{reasoning}}$ $=\text{VLM}(\widetilde{c}_{i},\texttt{Memory}^{l\text{-steps}}_{\text{visual}},\texttt{Prompt}_{\text{reasoning}})$ (5)
@@ 3.3 Tool-Augmented Self-Reflection
Given an Insufficient Reason from the current reasoning step, PANDA first queries its long chain-of-memory (Long-CoM, will be introduced in 3.4) to retrieve the most similar history reflection cases:
$\texttt{Experience}_{\text{reflection}}=\texttt{RetrieveTop1}(\texttt{Insufficient Reason},\texttt{Long-CoM}).$ (6)
@@
$\;\;\texttt{Visual Enhancement Info}=\widehat{c}_{j}\cup\{c_{k}\}\}.$
Here, text enhancement information includes summaries from tool outputs (e.g., detected objects, web search results), while visual enhancement information includes processed video clip $\widehat{c}_{j}$ and retrieved historical keyframes $\{c_{k}\}$.
@@
[…]
@@
$\texttt{Result}^{\text{reflection}}_{\text{reasoning}}=\text{VLM}(\widehat{c}_{j}\cup\{c_{k}\},\ \texttt{Memory}^{l\text{-steps}}_{\text{visual}},\ \texttt{Prompt}^{\text{refined}}_{\text{reasoning}}).$ (9)
If the returned status is Normal or Abnormal, PANDA resumes reasoning at the next timestep. If the status remains Insufficient, reflection is re-triggered. To prevent infinite loops, we limit the number of reflection rounds to $r$. If after $r$ rounds the result is still Insufficient, PANDA assigns a default anomaly score corresponding to the "Insufficient" status and skips the current segment and continues next timestep.
@@ 4.1 Experiment Setup
We evaluate PANDA on four benchmarks: UCF-Crime sultani2018real-22 , XD-Violence wu2020not-24 , UBnormal acsintoae2022ubnormal , and CSAD, which represent three distinct settings—multi-scenario, open-set, and complex scenario. UCF-Crime is a large-scale dataset comprising 1,900 long, untrimmed real-world surveillance videos. It covers 13 types of abnormal events such as fighting, abuse, stealing, arson, robbery, and traffic accidents. The training set includes 800 normal and 810 abnormal videos, while the test set consists of 150 normal and 140 abnormal videos. XD-Violence is another large-scale dataset focused on violence detection. It contains 4,754 videos collected from surveillance video, movies, and CCTV sources, encompassing 6 categories of anomaly events. The training and test sets include 3,954 and 800 videos, respectively. UBnormal is a synthetic open-set video anomaly detection dataset with a total of 543 videos. It defines 7 categories of normal events and 22 types of anomalies. Notably, 12 anomaly categories in the test set are unseen during training, making it a challenging benchmark for evaluating generalization under open-set conditions. CSAD is a complex-scene anomaly detection benchmark constructed in this work. It consists of 100 videos (50 normal and 50 abnormal), sampled from UCF-Crime, XD-Violence, and UBnormal. CSAD includes videos with challenging conditions such as low resolution, poor illumination, high noise levels, and long-range temporal anomalies. It is designed to assess model robustness in complex and degraded environments.
@@
We adopt Langgraph langchain2025langgraph to build the whole agent framework and all experiments are implemented using PyTorch imambi2021pytorch on the A6000 GPU. We use Qwen2.5VL-7B bai2025qwen2 as the VLM for perception and reasoning stages, and Gemini 2.0 Flash team2023gemini as the MLLM for planning and reflection. To improve the inference efficiency, the input video is 1FPS sampled, and one video clip containing $s=5$ frames is inferred at each time step. PANDA supports both offline and online reasoning modes. In offline reasoning mode, the perception phase is sampling $M=300$ frames uniformly for the whole video, while only the initial $M=10$ frames are sampled in online mode. The number of knowledge entries for each type of anomalous event in the anomaly knowledge base is $H=20$. The maximum number of reflection rounds $r$ is set to 3. The short CoM length $l=5$ during the reasoning stage. We retrieve the top $k=5$ anomaly rules from the anomaly knowledge base for each user query. More implementation details, including prompt templates and tool usage, are provided in the supplementary material.
Table 1: Comparisons with previous state-of-the-art methods on different datasets. "Expl." stands for "Explanation", indicating whether the output results include interpretations of the detected anomalies. Methods categorized as "Semi", "Weak", or "Instru-Tuned" require training data to adapt to specific scenarios or anomaly types. As a result, their performance degrades significantly when applied directly to unseen open-set scenarios. On the other hand, training-free specialized methods rely heavily on manual engineering to accommodate new scenes or anomalies, which limits their generalization capability.
@@
Table 2: We performed key hyperparameter analyses on the UCF-Crime dataset.
@@
4.2 Comparison with State-of-the-Art Methods
4.3 Analytic Results
Analysis of reflection round $r$.
@@ 4.3 Analytic Results
Analysis of rules number $k$.
Analysis of short CoM length $l$.
@@ 5 Conclusion
In this work, we presented PANDA, an agentic AI engineer for generalized VAD that eliminates the need for training data or manually crafted pipelines when faced with various real-world scenarios. PANDA integrates four core capabilities: self-adaptive scene-aware strategy planning, goal-driven heuristic reasoning, tool-augmented self-reflection, and a self-improving chain-of-memory. These modules work in concert to enable PANDA to adaptively detect anomalies across diverse, dynamic, and previously unseen environments. Our extensive experiments across multiple benchmarks, including multi-scene, open-set, and complex scenarios, validate PANDA’s strong generalization ability and robust performance without any training. These findings highlight PANDA’s potential as a generalist VAD solution for real-world scenes.
@@ C.2 Impact of Different MLLMs
    Input Clip Length UCF-Crime
@@
(b) Effect of input clip length.
@@ Appendix D Toolset Details
PANDA employs the YOLOWorld cheng2024yolo model pretrained on a wide set of open-world concepts. It supports fine-grained category-specific detection including actions like person hitting another”, person setting something on fire”, and “person stealing”. This enables robust anomaly-related scene understanding through bounding-box localization and category labels.
@@
PANDA integrates the Real-ESRGAN wang2021real model for resolution enhancement. It improves detail preservation and restores textures in low-resolution videos using a deep RRDBNet-based super-resolution pipeline.
@@
We use CLIP-based radford2021learning visual-textual retrieval to match current queries (e.g., “robbery incident”) with previously seen keyframes. Cosine similarity between CLIP embeddings is used for scoring relevance.
@@
PANDA leverages the Tavily Search API tavily for querying web content related to unknown or uncertain anomalies. Search results are parsed into structured summaries that can be referenced in the reasoning process.

### v2

@@ 1 Introduction
Technically, the proposed PANDA is distinguished by the following aspects. (1) self-adaptive scene-aware strategy planning. Faced with a new scene or user-defined anomaly detection requirements, PANDA first conducts environment perception and understanding, then retrieves relevant anomaly rules from an anomaly knowledge database. Based on the environment context information, a scene-adaptive Retrieval Augmented Generation (RAG) mechanism is designed to construct tailored anomaly detection plans. (2) Goal-driven heuristic reasoning. PANDA injects task-specific prompts guided by latent anomaly cues, which steer the reasoning process toward more accurate and focused decision-making. (3) Tool-augmented self-reflection. PANDA iteratively assesses uncertainty and activates a suite of curated tools, such as object detection, image retrieve, or web search, to acquire additional information and resolve ambiguous decision-making. (4) Self-improving chain-of-memory. PANDA integrates historical experiences to justify current reasoning decisions or self-reflection. By progressively accumulating contextual cues across temporal spans, it enhances both the stability and accuracy of its decisions over time. Taken together, PANDA embodies an agentic AI engineer that proactively perceives diverse environments, formulates adaptive strategies, performs goal-driven reasoning, and progressively improves through tool-augmented reflection and a chain-of-memory mechanism, enabling generalizable video anomaly detection across complex real-world scenarios.
@@ 2 Related Work
These approaches rely on varying levels of annotated data and typically fall into three categories: semi-supervised VAD hasan2016learning-12 ; lu2013abnormal-18 ; liu2018future-13 ; yang2023video-17 ; yang2022dynamic-15 , weakly-supervised VAD yang2024text-21 ; wu2024vadclip-25 ; wu2020not-24 ; sultani2018real-22 ; wu2024weakly ; lian2025vlial , and instruction-tuned VAD du2024uncovering-29 ; tang2024hawk-30 ; zhang2024holmes-07 ; zhang2024holmes-32 ; yang2025assistpda . For example, Ristea et al. ristea2024self proposed an efficient anomaly detection model based on a lightweight masked autoencoder. Yang et al. yang2024text-21 introduced a text prompt-driven pseudo-labeling and self-training framework for weakly-supervised VAD. Zhang et al. zhang2024holmes-32 presented a model combining an anomaly-focused temporal sampler with an instruction-tuned MLLM to detect anomalies. While these training-dependent methods often perform well within the domain of the training data, they typically suffer from sharp performance degradation when deployed in out-of-distribution environments or faced with novel anomaly types. This limits their applicability in the open-world scenarios where anomalies are diverse, and context-sensitive.
@@ 3.1 Self-adaptive Scene-aware Strategy Planning
Given a user-defined detection query $\texttt{User}_{\text{query}}$ and an input video sequence $V=\{f_{1},f_{2},\dots,f_{N}\}$ containing $N$ frames, PANDA uniformly samples $M$ keyframes $F=\{f_{1},f_{2},\dots,f_{M}\}$ and constructs a perception prompt $\texttt{Prompt}_{\text{perception}}$ combining $F$ with the $\texttt{User}_{\text{query}}$. This prompt is fed to a VLM, which returns structured environmental information:
EnvInfo $=\text{VLM}(F,\texttt{Prompt}_{\text{perception}})$ (1)
$=\{\texttt{Scene~Overview},\texttt{Potential~Anomalies},$
$\;\;\texttt{Weather~Condition},\texttt{Video~Quality}\}$
Here, Scene Overview provides a high-level summary of the scene, including location type (e.g., street, shop, parking lot) and observed activities. Potential Anomalies refers to types of suspicious behaviors that may plausibly occur in the current scene context. Weather Condition captures attributes such as time of day (day/night) and weather (e.g., sunny, rainy). Video Quality summarizes resolution and clarity (e.g., low-resolution, blurred, noisy).
@@ 3.2 Goal-Driven Heuristic Reasoning
Under the guidance of the detection strategy plan constructed in subsection 3.1, PANDA performs goal-driven heuristic reasoning using a VLM. Given the $\texttt{User}_{\text{query}}$, a clip-level video segment $V_{\text{clip}}=\{c_{1},c_{2},\dots,c_{T}\}$ (each video clip $c_{t}$ contains $s$ video frames), and the $\texttt{Plan}_{\texttt{strategy}}$, PANDA first applies the preprocessing tools specified in $\texttt{Plan}_{\texttt{strategy}}$ to obtain an enhanced video clip:
@@
Promptreasoning = $\text{Memory}^{\text{$l$-steps}}_{\text{text}}$, Potential Anomalies, $\texttt{Rules}_{a}$, Heuristic Prompt, Enhancement and Reflection Info}.
The fields Potential Anomalies, $\texttt{Rules}_{a}$, and Heuristic Prompt are directly inherited from the planning stage. The Enhancement and Reflection Info field incorporates information produced during the self-reflection stage (To be described in subsection 3.3), including tool-based refinements and updated anomaly rules and heuristic prompts. To enhance temporal awareness, PANDA equips a short-term memory component $\texttt{Memory}^{l\text{-steps}}_{\text{text}}$, which records the past $l$ reasoning steps as textual memory. In addition to textual memory, PANDA also maintains a corresponding visual memory stream $\texttt{Memory}^{l\text{-steps}}_{\text{visual}}$, which stores visual frames aligned with the latest $l$ steps, allowing the model to access fine-grained visual cues during inference.
@@
$\texttt{Result}_{\text{reasoning}}$ $=\text{VLM}(\widetilde{c}_{t},\texttt{Memory}^{l\text{-steps}}_{\text{visual}},\texttt{Prompt}_{\text{reasoning}})$ (6)
@@ 3.3 Tool-Augmented Self-Reflection
Given an Insufficient Reason from the current reasoning step, PANDA first queries its long chain-of-memory (Long CoM, will be introduced in 3.4) to retrieve the most similar history reflection cases:
$\texttt{Experience}_{\text{reflection}}=\texttt{RetrieveTop1}(\texttt{Insufficient Reason},\texttt{Long CoM}).$ (7)
@@
Here, Insufficient Reason refers to the underlying cause of decision uncertainty, inferred by the MLLM in conjunction with VLM output and contextual information such as environmental cues and anomaly rules. Tools to Use specifies the names of tools used for information enhancement and their corresponding parameters. New Anomaly Rule and New Heuristic Prompt represent the updated anomaly rule and the reformulated heuristic prompt, respectively.
@@
$\;\;\texttt{Visual Enhancement Info}=\widehat{c}_{t}\cup c_{s}\}.$
Here, Text Enhancement Info includes summaries from tool outputs (e.g., detected objects, web search results), while Visual Enhancement Info includes processed video clip $\widehat{c}_{t}$ and retrieved historical keyframe set $c_{s}=\{f_{1},f_{2},...,f_{s}\}$.
@@
$\texttt{Prompt}^{\text{refined}}_{\text{reasoning}}$ $=\texttt{Prompt}_{\text{reasoning}}\cup\{$ (10)
$\;\;\texttt{Text Enhancement Info},\ \texttt{New Anomaly Rule},$
$\;\;\texttt{New Heuristic Prompt}\}$
@@
$\texttt{Result}^{\text{reflection}}_{\text{reasoning}}=\text{VLM}(\widehat{c}_{t}\cup c_{s},\ \texttt{Memory}^{l\text{-steps}}_{\text{visual}},\ \texttt{Prompt}^{\text{refined}}_{\text{reasoning}}).$ (11)
If the returned status is Normal or Abnormal, PANDA resumes reasoning at the next timestep. If the status remains Insufficient, reflection is re-triggered. To prevent infinite loops, we limit the number of reflection rounds to $r$. If after $r$ rounds the result is still Insufficient, PANDA assigns a default anomaly score corresponding to the Insufficient status and skips the current segment and continues next timestep.
@@ 4.1 Experiment Setup
We evaluate PANDA on four benchmarks: UCF-Crime sultani2018real-22 , XD-Violence wu2020not-24 , UBnormal acsintoae2022ubnormal , and CSAD, which represent three distinct settings—multi-scenario (UCF-Crime and XD-Violence), open-set (UBnormal), and complex scenario (CSAD). UCF-Crime is a large-scale dataset comprising 1,900 long, untrimmed real-world surveillance videos. It covers 13 types of abnormal events such as fighting, abuse, stealing, arson, robbery, and traffic accidents. The training set includes 800 normal and 810 abnormal videos, while the test set consists of 150 normal and 140 abnormal videos. XD-Violence is another large-scale dataset focused on violence detection. It contains 4,754 videos collected from surveillance video, movies, and CCTV sources, encompassing 6 categories of anomaly events. The training and test sets include 3,954 and 800 videos, respectively. UBnormal is a synthetic open-set video anomaly detection dataset with a total of 543 videos. It defines 7 categories of normal events and 22 types of anomalies. Notably, 12 anomaly categories in the test set are unseen during training, making it a challenging benchmark for evaluating generalization under open-set conditions. CSAD is a complex-scene anomaly detection benchmark constructed in this work. It consists of 100 videos (50 normal and 50 abnormal), sampled from UCF-Crime, XD-Violence, and UBnormal. CSAD includes videos with challenging conditions such as low resolution, poor illumination, high noise levels, and long-range temporal anomalies. It is designed to assess model robustness in complex and degraded environments.
@@
We adopt Langgraph langchain2025langgraph to build the whole agent framework and all experiments are implemented using PyTorch imambi2021pytorch on the A6000 GPU. We use Qwen2.5VL-7B bai2025qwen2 as the VLM for perception and reasoning stages, and Gemini 2.0 Flash team2023gemini as the MLLM for planning and reflection. During the RAG process, the anomaly knowledge base and environment information are encoded via the all-MiniLM-L6-v2 model reimers-2020-multilingual-sentence-bert , with the knowledge base indexed using FAISS for efficient similarity retrieval. To improve the inference efficiency, the input video is sampled at 1 FPS, and a video clip of $s=5$ frames is inferred at each time step. PANDA supports both offline and online reasoning modes. In offline reasoning mode, the perception phase is sampling $M=300$ frames uniformly for the whole video, while only the initial $M=10$ frames are sampled in online mode. The number of knowledge entries for each type of anomalous event in the anomaly knowledge base is $H=20$. The maximum number of reflection rounds $r$ is set to 3. The short CoM length $l=5$ during the reasoning stage. We retrieve the top $k=5$ anomaly rules from the anomaly knowledge base for each user query. More implementation details, including prompt templates and tool usage, are provided in the supplementary material.
Table 1: Comparisons with previous state-of-the-art methods on different datasets. "Expl." stands for "Explanation", indicating whether the output results include interpretations of the detected anomalies. Methods categorized as "Semi", "Weak", or "Instru-Tuned" require training data to adapt to specific scenarios or anomaly types.
@@
4.2 Comparison with State-of-the-Art Methods
4.3 Analytic Results
Analysis of reflection round $r$.
Analysis of rules number $k$.
Analysis of short CoM length $l$.
Table 2: Key hyperparameter analyses on the UCF-Crime dataset.
@@ 5 Conclusion
In this work, we presented PANDA, an agentic AI engineer for generalized VAD that eliminates the need for training data or manually crafted pipelines when faced with various real-world scenarios. PANDA integrates four core capabilities: self-adaptive scene-aware strategy planning, goal-driven heuristic reasoning, tool-augmented self-reflection, and the self-improving chain-of-memory. These capabilities work in concert to enable PANDA to adaptively detect anomalies across diverse, dynamic, and previously unseen environments. Our extensive experiments across multiple benchmarks, including multi-scene, open-set, and complex scenarios, validate PANDA’s strong generalization ability and robust performance without any training. These findings highlight PANDA’s potential as a generalist VAD solution for real-world scenes.
@@ C.2 Impact of Different MLLMs
    Input Clip Length $s$ UCF-Crime
@@
(b) Effect of input clip length $s$.
@@ Appendix D Toolset Details
PANDA employs the YOLOWorld [42] model pretrained on a wide set of open-world concepts. It supports fine-grained category-specific detection including actions like “person hitting another”, “person setting something on fire”, and “person stealing”. This enables robust anomaly-related scene understanding through bounding-box localization and category labels.
@@
PANDA integrates the Real-ESRGAN [43] model for resolution enhancement. It improves detail preservation and restores textures in low-resolution videos using a deep RRDBNet-based super-resolution pipeline.
@@
PANDA uses CLIP-based [44] visual-textual retrieval to match current queries (e.g., “robbery incident”) with previously seen keyframes. Cosine similarity between CLIP embeddings is used for scoring relevance.
@@
PANDA leverages the Tavily Search API [45] for querying web content related to unknown or uncertain anomalies. Search results are parsed into structured summaries that can be referenced in the reasoning process.